• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name eschulte

    Log entry time 16:01:30 on June 18, 2007

    Entry number 205856

    keyword=Day Shift Log

    08:00 Shift begins taking data with 6% radiator in, Run #1327

    08:40 Mike Sealey from the target group has come over to contemplate the
    falling pressure in the 3He system. It may be a true leak, or radiation
    damage to transducer. Neither of these sound good. Target group has
    requested an access. I'll set it up with MCC & RC. The access is
    scheduled for 09:30.

    09:20 MCC let us know that they adjusted the "Box Supply".

    09:25 MCC called to get the last details of our access set up. I let
    them know that M. Sealey just needs to check the gas panel for now.

    09:30 MCC called to let us know that they were going to take the beam
    away and the ARM is on his way to the counting house.

    09:36 M. Sealey & ARM go down to hall. Paged RC to see if we can answer
    some of M. Sealey's questions about radiation levels in the hall.

    09:45 M. Sealey reported that the pressure change appeared to be real;
    however, the gas seems to have returned to the tank. This is good! He's
    re-pressurizing the cell. It is not clear what caused the gas to leak
    back into the tank (but fortunately it did). Nevertheless, the target
    pressure was low over the weekend.

    10:10 Hall back in beam permit. Requested 30 uA to continue running with
    radiator in (6%) Run #1329.

    10:30 It was noticed that run numbers 1320 through 1329 are labeled with
    the incorrect radiator thickness. They are labeled 5% although they
    should be 6%. All radiator in runs are taken with a 6% as far as I know.

    11:15 Change to radiator out.

    11:17 Requested 50uA to resume running with run #1330

    12:19 Change to radiator in.

    12:21 Requested 30uA to resume data taking with run #1331

    12:51 HRS-L Q3 has now take up drifting about (and Q2 has settled down a
    bit???).

    13:06 What sounded like a UPS alarm briefly went off. It stopped on its
    own.

    13:14 So now the HRS-L Q2/3 are going back and forth beeping. Q3 does,
    indeed seem to be drifing a bit low (effected run #1331). But it seems
    to be self-correcting.

    13:25 HRS-L Q3 drifted low again (2.7803) and re-regulated itself without
    intervention from the crew.

    13:30 Stopped run 1332 because we lost Q1 (HRS-R). We are attempting to
    reset it, however we were unsuccessful. Called M. Stevens (on-call) to
    get some assistance.

    13:41 Requested an access to hard reset the HRS-R Q1 power supply. On
    the upside, the HRS-L magnets seem to have settled down.

    13:59 Now the HRS-L Q3 has been drifting too high for the last several
    minutes.

    14:07 Q1 (HRS-R) Cryo Fault cleared mysteriously and is now resetting
    itself. HRS-L Q3 is still drifting about a bit.

    14:10 Request return to beam permit. HRS-R Q1 seems to be resetting
    normally.

    14:15 HRS-R Q1 successfully reset.

    14:17 Request 30uA after hall returned to Beam Permit. Resume running
    with run #1333.

    14:27 HRS-L Q3 drifted low again (run 1333). This is has been stable for
    several hours before this near constant instability occured.

    15:59 HRS-L Q3 has stablized for a little while, at least.